Can you please show me the graph of x+y=2, 2x-y=-8? I just don't get the graph, plot, slope stuff.
Answer by mananth(16946)

1 x + 1 y = 2
y=-x+2.............1
2 x -1 y = -8
y=2x+8 .............2
equation1
when x= 0, y= 2 ( 0 , 2 )
when y =0, x= 2 ( 2 , 0 )
Equation 2
when x= 0, y= 8 ( 0 , 8 )
when y =0, x= -4 ( -4 , 0 )

Red : y=-x+2
Green: y=2x+8